OpenStudy (anonymous):

A video game has been developed whereby the player must earn a certain number of experience points before progressing to the next level. On level 1, a player must earn 1 experience point. On level 2, a player must earn 6 experience points. On level 3, a player must earn 13 experience points. On level 4, a player must earn 22 experience points. Which recursive equation represents the pattern? an = an – 1 + n an = an – 1 + (4n + 1) <-- My answer (Am I correct?) an = an – 1 + (2n + 1) an = an – 1 + 3n

hero (hero):

So we have two sets n = {1,2,3,4...} a = {1,6,13,22...} As n increases, the number of experience points increases successively by i = {5, 7, 9, 11...} \[a_1 + 5 \\= 1 + 5 \\= { a_{(2 - 1)} + 2(2) + 1} \\ = 6 \\= a_2\] \[{a_3 = a_{(3 - 1)} + 2(3) + 1} \\ = a_2 + 7 \\= 6 + 7 \\= 13\]

hero (hero):

The answer is not what you chose @AdrianK

OpenStudy (anonymous):

i am not sure if i am reading it right ... is it A_n -1 or A_(n-1)

OpenStudy (anonymous):

So, it would be C?

hero (hero):

Yes, C is correct.

OpenStudy (anonymous):

Thanks again!! You are a life saver!

hero (hero):

yw